a linear function that passes through the points (-4,2) and (5,-16) what is the y intercept
Answer by dkppathak(439)

a linear function that passes through the points (-4,2) and (5,-16) what is the y intercept
slope (m)=y2-y1/x2-x1
m=-16-2/5-(-4)=-18/9
linear function
y-y1=m(x-x1)
y-2=-18/9(x+4)
9(y-2)=-18(x+4)
9y-18=-18x-72
9y=-18x-72+18
9y=-18x-54 compression y=mx+c where c is intercept on y axis
so y intercept is -54